We've still be experiencing 20 knot winds on the ground at Halley and not great conditions at SANAE either. The weather is predicted to improve at Halley through the evening and into tomorrow so we are hoping for a launch opportunity Sunday. We lost contact with Payload 2C this morning at 7:46:33 UTC. The telemetry throughput for this payload has been lower on average than the others (~85%) so we suspect a modem issue may be responsible. The payload will terminate itself after 38 hours without contact but we will continue to attempt contact until that point. On a more positive note, this is the first hardware problem we've had this year and we have now satisfied our comprehensive success criteria in terms of the number of hours of observations. The attached plot shows a histogram of the total number of hours of observation at various L-values. To date, we have acquired 1380 hours of observations between L=3-7 (based on T89 field model with Kp=2).
Happy Burn's Night!